# Anti-Leibniz algebras: A non-commutative version of mock-Lie algebras

### v1

@@ 1 Introduction and preliminary results
Anti-ass trialg.Theorem 4.10Anti-Leibniz trialg.Anti-ass. dialg.Theorem2.4 Anti-Leibniz alg. Anti-ass. alg.Prop.3.10Proposition 4.9Anti-Comm.Mock-Lie alg.Theorem3.6Theorem 4.3
@@
It is known that any anti-associative algebra $(A,\star)$ gives rise to a a mock-Lie algebra $(A,\circ)$, via anti-commutator, that is $x\circ y:=x\star y+y\star x,\;\forall x,y\in A$.
@@
Let $(A,\star)$ be an anti-associative algebra, $V$ be a vector space and $l,r:A\to End(V)$ are linear maps. Then $(V,\varrho,\mu)$ is a representation of $A$ if and only if $(A\oplus V,\star_{A\oplus V})$ is an anti-associative algebra, where
@@ 2.1 Definition and Examples
If $(A,\{\cdot,\cdot\})$ is a anti-Leibniz algebra then $(A,\{\cdot,\cdot\}\tau_{12})$ is a right anti-Leibniz algebra.
@@ 2.2 Representations of anti-Leibniz algebras
A representation of an anti-Leibniz algebra $(A,\{\cdot,\cdot\})$ is a triple $(V;\mathfrak{l},\mathfrak{r})$, where $V$ is a vector space, $\mathfrak{l},\mathfrak{r}:A\to End(V)$ are linear maps satisfying:
@@
Let $(A,\{\cdot,\cdot\})$ be an anti-Leibniz algebra. Then, using Definitions 2.1 and 2.3, we have for any $x,y,z\in A$ and $u,v,w\in V$
@@
Let $(V;\mathfrak{l},\mathfrak{r})$ be a representation of an anti-Leibniz algebra $(A,\{\cdot,\cdot\})$. Define the two linear maps $\mathfrak{l}^{*},\mathfrak{r}^{*}:A\longrightarrow End(V^{*})$, respectively by
@@
Under the above notations, $\big{(}V^{*};\mathfrak{l}^{*},\mathfrak{l}^{*}-\mathfrak{r}^{*}\big{)}$ is a representation of $(A,\{\cdot,\cdot\})$, which is called the dual representation of $(V;\mathfrak{l},\mathfrak{r})$.
@@
By (2.7), for all $x,y\in A,~{}v\in V$ and $\chi\in V^{*}$, we have
$\langle\mathfrak{l}^{*}(\{x,y\})\chi,v\rangle$ $=$ $\langle\chi,\mathfrak{l}(\{x,y\})v\rangle$
$=$ $-\langle\chi,\mathfrak{l}(x)\mathfrak{l}(y)v+\mathfrak{l}(y)\mathfrak{l}(x)v\rangle$
$=$ $-\langle\mathfrak{l}^{*}(y)\mathfrak{l}^{*}(x)\chi,v\rangle-\langle\mathfrak{l}^{*}(x)\mathfrak{l}^{*}(y)\chi,v\rangle$
$=$ $\langle-\mathfrak{l}^{*}(y)\mathfrak{l}^{*}(x)\chi-\mathfrak{l}^{*}(x)\mathfrak{l}^{*}(y)\chi,v\rangle.$
@@
$\langle\big{(}\mathfrak{l}^{*}(\{x,y\})-\mathfrak{r}^{*}(\{x,y\})\big{)}\chi,v\rangle$
$=$ $\langle\chi,\mathfrak{l}(\{x,y\})v-\mathfrak{r}(\{x,y\})v\rangle$
$=$ $\langle\chi,-\mathfrak{l}(x)\mathfrak{l}(y)v-\mathfrak{l}(y)\mathfrak{l}(x)v+\mathfrak{l}(x)\mathfrak{r}(y)v+\mathfrak{r}(y)\mathfrak{l}(x)v\rangle$
$=$ $-\langle\mathfrak{l}^{*}(y)\mathfrak{l}^{*}(x)\chi,v\rangle-\mathfrak{l}^{*}(x)\mathfrak{l}^{*}(y)\chi,v\rangle$
$+\langle\mathfrak{r}^{*}(y)\mathfrak{l}^{*}(x)\chi,v\rangle+\langle\mathfrak{l}^{*}(x)\mathfrak{r}^{*}(y)\chi,v\rangle$
$=$ $-\langle\mathfrak{l}^{*}(x)(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))\chi,v\rangle-\langle(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))\mathfrak{l}^{*}(x)\chi,v\rangle$
$=$ $\langle-\mathfrak{l}^{*}(x)(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))\chi-(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))\mathfrak{l}^{*}(x)\chi,v\rangle.$
@@
$\langle(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))(\mathfrak{l}^{*}(x)\chi),v\rangle$
$=$ $\langle\chi,(\mathfrak{l}(x)\mathfrak{l}(y)-\mathfrak{l}(x)\mathfrak{r}(y))v\rangle$
$=$ $\langle\chi,(\mathfrak{l}(x)\mathfrak{l}(y)-\mathfrak{r}(x)\mathfrak{l}(y)-\mathfrak{l}(x)\mathfrak{r}(y)+\mathfrak{r}(x)\mathfrak{r}(y))v\rangle$
$=$ $\langle\chi,(\mathfrak{l}(x)-\mathfrak{r}(x))(\mathfrak{l}(y)-\mathfrak{r}(y))v\rangle$
$=$ $\langle(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))(\mathfrak{l}^{*}(x)-\mathfrak{r}^{*}(x))\chi,v\rangle$
for all $x,y\in A,v\in V$ and $\chi\in V^{*}$. Thus, we have
@@
Therefore $\big{(}V^{*};\mathfrak{l}^{*},\mathfrak{l}^{*}-\mathfrak{r}^{*}\big{)}$ is a representation of $(A,\{\cdot,\cdot\})$. ∎
@@
Under the above notations, $\big{(}A^{*};L^{*},L^{*}-R^{*}\big{)}$ is a representation of $(A,\{\cdot,\cdot\})$, which is called the coadjoint representation of $(A;L,R)$.
@@ 3 Embedding tensor on mock-Lie algebras and anti-Leibniz algebras
Let $\mathcal{K}:V\rightarrow A$ be an embedding tensor on an anti-associative algebra $(A,\star)$ with respect to a representation $(V,\varrho,\mu)$. Then, $V$ carries an anti-associative dialgebra structure given by
@@ 4 Anti-Leibniz trialgebras
With the above notation, $(A\oplus B,\{\cdot,\cdot\}_{A\oplus B},\circ_{A\oplus B})$ is a anti-Leibniz trialgebra, called the hemisemidirect product anti-Leibniz trialgebra and denoted by $A\oplus_{ALeibT}B$.
@@
$(a\triangleleft_{\mathcal{H}}b)\triangleleft_{\mathcal{H}}c+a\triangleleft_{\mathcal{H}}(b\triangle_{\mathcal{H}}c)$ $=$ $\mu({\mathcal{H}}(c))(\mu({\mathcal{H}}b)a)+\mu({\mathcal{H}}(b\star_{B}c))a$

### v2

@@ 1 Introduction and preliminary results
Anti-ass trialg.Theorem 4.12Anti-Leibniz trialg.Anti-ass. dialg.Theorem2.5 Anti-Leibniz alg. Anti-ass. alg.Prop.3.12Proposition 4.11Anti-Comm.Mock-Lie alg.Theorem3.6Theorem 4.3
@@
It is known that any anti-associative algebra $(A,\star)$ gives rise to a mock-Lie algebra $(A,\circ)$, via anti-commutator, that is
$x\circ y:=x\star y+y\star x,\;\forall\ x,y\in A.$ (1.7)
@@
Example 1.3.
Let $(A,\star)$ be an anti-associative algebra and define the linear maps $l,r:A\to End(A)$ by $l(x)(y)=r(y)(x)=x\star y$. Then $(A,l,r)$ is a representation of $A$ on itself, called the adjoint representation.
@@
Let $(A,\star)$ be an anti-associative algebra, $V$ be a vector space and $\varrho,\mu:A\to End(V)$ are linear maps. Then $(V,\varrho,\mu)$ is a representation of $A$ if and only if $(A\oplus V,\star_{A\oplus V})$ is an anti-associative algebra, where
@@ 2.1 Definition and Examples
Definition 2.2.
Let $(A,\{\cdot,\cdot\})$ be an anti-Leibniz algebra and $I$ be a sub-vector space of $A$.
  1. 1.
    $I$ is called a subalgebra of $A$ if $\{I,I\}\subseteq I$.
  2. 2.
    $I$ is called an ideal of $A$ if $\{I,A\}\subseteq I$.
Example 2.2.
Let $(A,\{\cdot,\cdot\})$ be an anti-Leibniz. The sub-vector space $I_{A}$ spanned by
$\bigg{\{}\{x,y\}-\{y,x\},\;x,y\in A\bigg{\}}$
is an ideal of $A$; in fact $\{I_{A},A\}=\{0\}$. It is clear that $A$ is a mock-Lie algebra if and only if $I_{A}=\{0\}$. Therefore, the quotient algebra $A/I_{A}$ is a mock-Lie algebra. The space $I_{A}$ is called the anti-Leibniz kernel of $A$.
@@
If $(A,\{\cdot,\cdot\})$ is a anti-Leibniz algebra, then $(A,\{\cdot,\cdot\}^{op})$ is a right anti-Leibniz algebra, where $\{x,y\}^{op}=\{y,x\}$, for all $x,y\in A$.
@@ 2.2 Representations of anti-Leibniz algebras
A representation of an anti-Leibniz algebra $(A,\{\cdot,\cdot\})$ is a triple $(V,\mathfrak{l},\mathfrak{r})$, where $V$ is a vector space, $\mathfrak{l},\mathfrak{r}:A\to End(V)$ are linear maps satisfying:
@@
Let $(A,\{\cdot,\cdot\})$ be an anti-Leibniz algebra. Then, using Definitions 2.1 and 2.4, we have for any $x,y,z\in A$ and $u,v,w\in V$
@@
Let $(V,\mathfrak{l},\mathfrak{r})$ be a representation of an anti-Leibniz algebra $(A,\{\cdot,\cdot\})$. Define the two linear maps $\mathfrak{l}^{*},\mathfrak{r}^{*}:A\longrightarrow End(V^{*})$, respectively by
@@
Under the above notations, $\big{(}V^{*},\mathfrak{l}^{*},\mathfrak{l}^{*}-\mathfrak{r}^{*}\big{)}$ is a representation of $(A,\{\cdot,\cdot\})$, which is called the dual representation of $(V,\mathfrak{l},\mathfrak{r})$.
@@
By (2.7), for all $x,y\in A,~{}v\in V$ and $\xi\in V^{*}$, we have
$\langle\mathfrak{l}^{*}(\{x,y\})\xi,v\rangle$ $=$ $\langle\xi,\mathfrak{l}(\{x,y\})v\rangle$
$=$ $-\langle\xi,\mathfrak{l}(x)\mathfrak{l}(y)v+\mathfrak{l}(y)\mathfrak{l}(x)v\rangle$
$=$ $-\langle\mathfrak{l}^{*}(y)\mathfrak{l}^{*}(x)\xi,v\rangle-\langle\mathfrak{l}^{*}(x)\mathfrak{l}^{*}(y)\xi,v\rangle$
$=$ $\langle-\mathfrak{l}^{*}(y)\mathfrak{l}^{*}(x)\xi-\mathfrak{l}^{*}(x)\mathfrak{l}^{*}(y)\xi,v\rangle.$
@@
$\langle\big{(}\mathfrak{l}^{*}(\{x,y\})-\mathfrak{r}^{*}(\{x,y\})\big{)}\xi,v\rangle$
$=$ $\langle\xi,\mathfrak{l}(\{x,y\})v-\mathfrak{r}(\{x,y\})v\rangle$
$=$ $\langle\xi,-\mathfrak{l}(x)\mathfrak{l}(y)v-\mathfrak{l}(y)\mathfrak{l}(x)v+\mathfrak{l}(x)\mathfrak{r}(y)v+\mathfrak{r}(y)\mathfrak{l}(x)v\rangle$
$=$ $-\langle\mathfrak{l}^{*}(y)\mathfrak{l}^{*}(x)\xi,v\rangle-\mathfrak{l}^{*}(x)\mathfrak{l}^{*}(y)\xi,v\rangle$
$+\langle\mathfrak{r}^{*}(y)\mathfrak{l}^{*}(x)\xi,v\rangle+\langle\mathfrak{l}^{*}(x)\mathfrak{r}^{*}(y)\xi,v\rangle$
$=$ $-\langle\mathfrak{l}^{*}(x)(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))\xi,v\rangle-\langle(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))\mathfrak{l}^{*}(x)\xi,v\rangle$
$=$ $\langle-\mathfrak{l}^{*}(x)(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))\xi-(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))\mathfrak{l}^{*}(x)\xi,v\rangle.$
@@
$\langle(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))(\mathfrak{l}^{*}(x)\xi),v\rangle$
$=$ $\langle\xi,(\mathfrak{l}(x)\mathfrak{l}(y)-\mathfrak{l}(x)\mathfrak{r}(y))v\rangle$
$=$ $\langle\xi,(\mathfrak{l}(x)\mathfrak{l}(y)-\mathfrak{r}(x)\mathfrak{l}(y)-\mathfrak{l}(x)\mathfrak{r}(y)+\mathfrak{r}(x)\mathfrak{r}(y))v\rangle$
$=$ $\langle\xi,(\mathfrak{l}(x)-\mathfrak{r}(x))(\mathfrak{l}(y)-\mathfrak{r}(y))v\rangle$
$=$ $\langle(\mathfrak{l}^{*}(y)-\mathfrak{r}^{*}(y))(\mathfrak{l}^{*}(x)-\mathfrak{r}^{*}(x))\xi,v\rangle$
for all $x,y\in A,v\in V$ and $\xi\in V^{*}$. Thus, we have
@@
Therefore $\big{(}V^{*},\mathfrak{l}^{*},\mathfrak{l}^{*}-\mathfrak{r}^{*}\big{)}$ is a representation of $(A,\{\cdot,\cdot\})$. ∎
@@
Under the above notations, $\big{(}A^{*},L^{*},L^{*}-R^{*}\big{)}$ is a representation of $(A,\{\cdot,\cdot\})$, which is called the coadjoint representation of $(A;L,R)$.
@@ 3 Embedding tensor on mock-Lie algebras and anti-Leibniz algebras
The embedding tensor of an anti-associative algebra $A$ with respect to the adjoint representation $(A,l,r)$ is called an averaging operator. In this case the identity (3.5) can be written as
$\mathcal{K}(x)\star\mathcal{K}(y)=\mathcal{K}(\mathcal{K}(x)\star y)=\mathcal{K}(x\star\mathcal{K}(y)),\,\forall\,x,y\in A.$
@@
Let $(V,\varrho,\mu)$ be a representation of an anti-associative algebra $(A,\star)$. Then $(V,\pi=\varrho+\mu)$ is a representation of the induced mock-Lie algebra $(A,\circ)$ given in (1.7). Moreover, if $\mathcal{K}:V\to A$ is an embedding tensor on the anti-associative algebra $(A,\star)$ with respect to the representation $(V,\varrho,\mu)$. Then $\mathcal{K}$ is an embedding tensor on the mock-Lie algebra $(A,\circ)$ with respect to the representation $(V,\pi)$.
Proof.
For any $x,y\in A$ and $u\in V$, by (1.11) we have
$\pi(x\circ y)u+\pi(x)(\pi(y)u)+\pi(y)(\pi(x)u)$
$=$ $(\varrho+\mu)(x\star y+y\star x)u+(\varrho+\mu)(x)((\varrho+\mu)(y)u)+(\varrho+\mu)(y)((\varrho+\mu)(x)u)$
$=$ $\varrho(x\star y)u+\varrho(y\star x)u+\mu(x\star y)u+\mu(y\star x)u+\varrho(x)\varrho(y)u+\varrho(x)\mu(y)u+\mu(x)\varrho(y)u+\mu(x)\mu(y)u$
$+\varrho(y)\varrho(x)u+\varrho(y)\mu(x)u+\mu(y)\varrho(x)u+\mu(y)\mu(x)u$
$=$ $0$
Thus $(V,\pi=\varrho+\mu)$ is a representation of $(A,\circ)$. On the other hand, if $\mathcal{K}$ is an embedding tensor on the anti-associative algebra $(A,\star)$, then for any $u,v\in V$, we have
$\mathcal{K}(u)\circ\mathcal{K}(v)$ $=\mathcal{K}(u)\star\mathcal{K}(v)+\mathcal{K}(v)\star\mathcal{K}(u)$
$=\mathcal{K}(\varrho(\mathcal{K}(u))v)+\mathcal{K}(\mu(\mathcal{K}(u))v)$
$=\mathcal{K}(\pi(\mathcal{K}(u))v).$
Hence, $\mathcal{K}$ is also an embedding tensor on the induced mock-Lie algebra $(A,\circ)$ with respect to $(V,\pi)$. ∎
Corollary 3.11.
Any averaging operator $\mathcal{K}:A\to A$ on an anti-associative algebra $(A,\star)$ is an averaging operator on the induced mock-Lie algebra $(A,\circ)$.
Proposition 3.12.
Let $\mathcal{K}:V\rightarrow A$ be an embedding tensor on an anti-associative algebra $(A,\star)$ with respect to the representation $(V,\varrho,\mu)$. Then, $V$ carries an anti-associative dialgebra structure given by
@@ 4 Anti-Leibniz trialgebras
A homomorphic embedding tensor on a mock Lie algebra $A$ with respect to the adjoint representation $(A,L)$ is called a homomorphic averaging operator.
@@
With the above notation, $(A\oplus B,\{\cdot,\cdot\}_{A\oplus B},\circ_{A\oplus B})$ is an anti-Leibniz trialgebra, called the hemisemidirect product anti-Leibniz trialgebra and denoted by $A\oplus_{ALeibT}B$.
@@
A homomorphic embedding tensor on the anti-associative algebra $(A,\star_{A})$ with respect to the adjoint representation $(A,l,r)$ is called homomorphic averaging operator.
Proposition 4.9.
Let $(B,\varrho,\mu,\star_{B})$ be an action of an anti-associative algebra $(A,\star_{A})$. Then $(B,\pi=\varrho+\mu,\circ_{B})$ is an action of the induced mock Lie algebra $(A,\circ_{A})$, where "$\circ_{A}"$ and "$\circ_{B}"$ are the anti-commutator given in (1.7). Moreover, if $\mathcal{H}:B\to A$ is a homomorphic embedding tensor on the anti-associative algebra $(A,\star_{A})$ with respect to the action $(B,\varrho,\mu,\star_{B})$. Then $\mathcal{H}$ is a homomorphic embedding tensor on the mock-Lie algebra $(A,\circ_{A})$ with respect to the action $(B,\pi,\circ_{B})$.
Proof.
According to Proposition 3.10 $(B,\pi)$ is a representation of $(A,\circ)$, it remains to check (4.1). For any $x\in A$ and $a,b\in B$, by (4.5) we have
$\pi(x)(a\circ_{B}b)+a\circ_{B}\pi(x)b+b\circ_{B}\pi(x)a$
$=$ $(\varrho+\mu)(x)(a\star_{B}b+b\star_{B}a)+a\star_{B}(\varrho+\mu)(x)b+(\varrho+\mu)(x)b\star_{B}a+b\star_{B}(\varrho+\mu)(x)a$
$+(\varrho+\mu)(x)a\star_{B}b$
$=$ $\varrho(x)(a\star_{B}b)+\varrho(x)(b\star_{B}a)+\mu(x)(a\star_{B}b)+\mu(x)(b\star_{B}a)+a\star_{B}(\varrho(x)b)+a\star_{B}(\mu(x)b)$
$+(\varrho(x)b)\star_{B}a+(\mu(x)b)\star_{B}a+b\star_{B}(\varrho(x)a)+b\star_{B}(\mu(x)a)+(\varrho(x)a)\star_{B}b+(\mu(x)a)\star_{B}b$
$=$ $0$
On the other hand, if $\mathcal{H}$ is a homomorphic embedding tensor on the anti-associative algebra $(A,\star_{A})$ with respect to $(B,\varrho,\mu,\star_{B})$, then according to Proposition 3.10 $\mathcal{H}$ is an embedding tensor on the mock Lie algebra $(A,\circ_{A})$ with respect to $(B,\pi)$. Now, it remains to verify that $\mathcal{H}$ is also a morphism. For any $a,b\in B$, we have
$\mathcal{H}(a\circ_{B}b)$ $=\mathcal{H}(a\star_{B}b)+\mathcal{H}(b\star_{B}a)$
$=\mathcal{H}(a)\star_{A}\mathcal{H}(b)+\mathcal{H}(b)\star_{A}\mathcal{H}(a)$
$=\mathcal{H}(a)\circ_{A}\mathcal{H}(b).$
Hence, the result follows. ∎
Corollary 4.10.
A homomorphic averaging operator $\mathcal{H}$ on a anti-associative algebra $(A,\star_{A})$ is a homomorphic averaging operator on the induced mock-Lie algebra $(A,\circ_{A})$.
@@
$(a\triangleleft_{\mathcal{H}}b)\triangleleft_{\mathcal{H}}c+a\triangleleft_{\mathcal{H}}(b\triangle_{\mathcal{H}}c)$ $=$ $\mu({\mathcal{H}}(c))(\mu({\mathcal{H}}(b))a)+\mu({\mathcal{H}}(b\star_{B}c))a$
@@ 5 Further discussion
Acknowledgement
The authors would like to thank the referee for valuable comments and suggestions on this article.